Marginal Units of Pesticide Runoff Abated Marginal Benefit of Abatement Marginal Cost of Abatement Farm A Cost of Abatement Farm B %20 $21 $23 $27 $31 %35 $38 $40 $45 $%55 $30 %90 $80 $75 $60 50 $40 $37 %35 %15 %10 $34 $37 $40 $45 $50 $55 $57 $%59 $65 2 3 4 5 6 7 8 10 uestion 14 The government proposes an emissions standard mandate for all farms to (uniformly) reduce pesticide runoff by 40% (Assume Farm A and Farm B both have total pesticide runoffs of 10 units.) A. What is the total benefit from this abatement policy? B. What is the total cost from this abatement policy?
Answer #1

Given the marginal benefits and marginal costs of abatement, and that both Farm A and B have 10 units of pesticide runoff, a reduction of 40% is equivalent to reducing pesticide runoff by 4 units each.

(a) Marginal benefit of abatement by Farm A = MB ( 4) = $ 60. Similarly, marginal benefit to farm B = $60

thus, total benefit of abating 4 units each = 60 (4) + 60 ( 4) = $480.

(b) Marginal cost of abatement of Farm A = $ 40 and marginal cost of abatement of farm B = $27

Thus, total cost of abating 4 units of pesticide runoff each = 40 (4) + 27 (4) = $268

Ans. Total benefit of abatement policy = $480 and total cost = $268